13.2 Electric Force. Coulombs Law Coulombs Lau Firi a F12= 41 o 2f12-Note that Fi2 is the force on l by 2; and 12-71-72 points from 2 to 1.fi 7 is the unit vector in the direction of In general for multiple di rete charges, net orce on charge 1, F (Pairwise vector sum; superposition principle). For convenience, ke = 9 x 109 appropriate units. Think How is N3L applicable here? Σ in 4
12. Two small, identical balls of mass m each are initially given charges of -2Q and +8Q. They experience a Coulomb force F when they are a distance d from each other. They are then brought into contact until the charge on them equilibrates. What is the Coulomb force exerted when they are separated again by a distance d, in terms of F
Answer #1

13.2)

Charge 0 is at rest located at vec{r}_1.

Charge Q_2 is at rest located at vec{r}_2.

r12 12T12 r12 = , r21 = r21r21 - , Y12-一7.21 , 1221

Force on charge 0 due to charge Q_2 is 1 0192 2T12 12 4TE0o 12

Force on charge Q_2 due to charge 0 is 1 0192 Fzi 21 21ー

1 012 4πε0 r12 1 012 4πε0 r21 F12 +F21

1 012 1 102 Fn+ F21 =

vec{F}_{12}+vec{F}_{21}=0

12 --「21

Force on charge 0 due to charge Q_2 is equal and opposite to force on charge Q_2 due to charge 0. (Newton's third law)

Hence Newton's third law is applicable here.

---------------------

12)

Initially the charges on two balls are -2Q,+8Q, separated by a distance d

Force of attraction between the two balls is F=rac{1}{4piepsilon_0}rac{2Q*8Q}{d^2}=rac{1}{4piepsilon_0}rac{16Q^2}{d^2}

The two balls brought in contact, and the charge on two balls are equal, charge on each ball is Q'=rac{-2Q+8Q}{2}=+3Q

Again the two balls are separated through a distance of d.

Force of repulsion between the two balls is F'=rac{1}{4piepsilon_0}rac{3Q*3Q}{d^2}=rac{1}{4piepsilon_0}rac{9Q^2}{d^2}=rac{9}{16}F

----------------------